Part 1: Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form.
Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

Answer by Alan3354(69443)

Part 1: Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form.
|+x +y +1|
|+3 +4 +1| = 0
|-3 -8 +1|
----------------
x*(4+8) - y*(3+3) + (-24+12) = 0
12x - 6y = 12
2x - y = 2
y = 2x - 2
--------------
Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.
What point? I used both.